# Learning Continually by Spectral Regularization

### v1

@@ 6 Conclusion
In this paper, we investigated the connection between initialization, trainability and regularization. We identified that controlling the singular values so that they do not deviate significantly from their values at initialization promotes diverse gradients for neural networks that sustains trainability throughout continual learning. As the first approach to preserve trainability properties induced by initialization, building on recent work, we proposed to regularize the distribution of individual parameters during training towards the distribution at initialization. We then proposed to use spectral regularization as a way of controlling the spectral norm of each layer, thereby sustaining the trainability properties that are present at initialization. Our experiments show that spectral regularization is more performant and less sensitive to hyperparameters than other regularizers across datasets, nonstationarities, and architectures. While several methods mitigate loss of trainability to a varying degree, learning continually with spectral regularization is robust, achieving high initial and sustained performance. Finally, we show that spectral regularization is also capable of improving generalization with both Vision Transformer and ResNet-18 on continual versions of SVHN2, CIFAR10 and CIFAR100.
@@ Appendix A Additional Details
A.1 Common Assumptions Regarding Tasks
We denote the parameters of a randomly initialized neural network trained only on task $\tau$ for $T$ iterations as $\theta_{\tau}^{(T)}=\theta^{(T-1)}_{\tau}-\alpha\nabla_{\theta}J_{\tau}(\theta)\big{|}_{\theta=\theta_{\tau}^{(t-1)}}$, with $\theta^{(0)}_{\tau}$ being sampled from a standard initialization distribution, such as Glorot [18] or He initialization [29]. The equal-difficulty assumption can be formalized by assuming that the objective value, $J_{\tau}(\theta_{\tau}^{(T)})$, is statistically independent of $\tau$. That is, while the data depends on the task, the objective value after training does not depend on the task. Note that the equal-difficulty assumption is key: otherwise it would be difficult to disambiguate whether a performance decrease is due to a more difficult task or due to loss of trainability in the learning algorithm.
A.2 Illustrative Example: Large Spectral Norm Can Impede Trainability
We show that, given a family of solutions on the first task, preference should be given to solutions that have a spectral norm that is closer to one. We consider 2 dimensional binary and orthogonal inputs, $x_{1}=[0,1]^{\top}$ and $x_{2}=[1,0]^{\top}$, with corresponding binary targets, $y_{1}=y_{2}=1$. We consider the squared loss, and use a mutli-layer perceptron with one hidden layer and two units, $f_{w_{2},w_{1}}(x)=w_{2}\texttt{ReLU}(w_{1}x)$. A common observation regarding solutions found for deep neural networks is that they are low rank [54, 55, 56, 57]. In addition, it has been observed that large outliers in the spectrum of a deep network emerge during training, meaning that the largest singular values of the weight matrices tend to be much larger than their smallest singular values [58, 22]. Consider one such solution where $w_{1}=\begin{bmatrix}-1&c\\ \nicefrac{{1}}{{c}}&-1\end{bmatrix}$, and $w_{2}=[\nicefrac{{1}}{{c}},c]\in\mathbb{R}^{1\times 2}$. The input-output function is invariant under different values of $c$, achieving zero error for any particular choice. However, the parameter matrix for the hidden activations, $w_{1}$, is low rank because it only has a single non-zero singular value that is not invariant. That is, the largest singular value of $w_{1}$ depends on the choice of $c$, $\|w_{1}\|_{2}:=\sigma_{max}(w_{1})=c+\nicefrac{{1}}{{c}}$. In addition, the gradients and parameter updates depend on $c$. Thus, if the targets were to change, $y_{1}^{\prime}=0$, the imbalance of singular values due to very large or small values of $c$ can impede trainability:
@@ A.2 Illustrative Example: Large Spectral Norm Can Impede Trainability
where $\ell_{1}=\ell(f_{w_{2},w_{1}}(x_{1}),y^{\prime}_{1})$, and $\alpha$ is a step-size. One of the consequences of the imbalanced singular values is that the per-example gradients are poorly conditioned [59]. That is, when we sample $x_{1}$, there will be a large gradient for $w_{2}$, $\|\nabla_{w_{2}}\ell(f_{w_{2},w_{1}}(x_{1}),y_{1}^{\prime})\|^{2}=c$ but a small gradient for $w_{1}$, $\|\nabla_{w_{1}}\ell(f_{w_{2},w_{1}}(x_{1}),y_{1}^{\prime})\|^{2}=\nicefrac{{1}}{{c}}$. Thus, a sufficiently small step-size that stabilizes the update to $w_{2}$ leads to slow learning on $w_{1}$. A similar result holds if the target $y_{2}$ were to change In Figure 7, we demonstrate that learning is more stable with a smaller spectral norm in this example. For higher dimensional weight matrices, we may expect that the imbalance of the singular values leads to low gradient diversity where particular per-example gradients give rise to outliers with a large gradient magnitudes. Thus, all else being equal, a solution with a smaller spectral norm ($c\approx 1$) is preferable for the purposes of continual learning.
@@
A.3 Differences Between Continual and Non-continual Learning
@@
A.4 Spectral Norm and Lipschitz Constant
One other potential reason for why a growing spectral norm is problematic is from the theory convex optimisation. For the ReLU activation function, the Lipschitz constant of the layer map, $h_{l+1}=\phi(\theta_{l}h_{l})$, is equal to the spectral norm of the weight matrix. We can thus bound the Lipschitz constant of the entire neural network above by the product of the layer-wise Lipschitz constants [60]. Putting these results together, we conclude that during continual learning, the Lipschitz constant of each layer is increasing and thus the Lipschitz constant of the entire network is also increasing. This is problematic from the standpoint of optimisation because gradient descent only converges on convex optimisation problems when the step-size is smaller than $\frac{2}{L}$, where $L$ is the Lipschitz constant. Thus, if the Lipschitz constant grows then, eventually, the chosen step-size will be too large and gradient descent will not converge locally on the task.
A.5 More Details on Regularization and Trainability
One other important advantage of regularization is that it has well understood effects for ensuring trainability in approaches outside of deep learning, such as degenerate linear regression and inverse problems [61].
@@ A.5 More Details on Regularization and Trainability
A.6 More Details on Wasserstein Regularizer
At initialization, the parameters at layer $l$ are sampled i.i.d. $\mathbf{\theta}_{i,j}\sim p^{(l,0)}(\theta)$ according to some pre-determined distribution, such as the Glorot initialization [18]. The distribution of parameters at iteration $t$ during training and for any particular layer, denoted by $p^{(l,t)}$, is no longer known and does not have a closed form. However, it is still possible to regularize the empirical distribution towards the initialization distribution by using the empirical Wasserstein metric [39]. We denote the flattened parameter matrix for layer $l$ at time $t$ by $\mathbf{\bar{\theta}}^{(l,t)}$. The squared Wasserstein-2 distance between the distribution of parameters at initialization and the current parameter distribution is defined as
$\mathcal{R}_{\tau(t)}(\theta,\theta^{(0)})=\sum_{i=1}^{d}\left(\bar{\theta}_{(i)}^{(l,t)}-\bar{\theta}_{(i)}^{(l,0)}\right)^{2}.$
Unlike the regenerative regularizer, the Wasserstein regularizer takes the difference of the order statistics. Thus, the regenerative regularizer is always larger because the Wasserstein regularizer takes the difference in the sorted values, $\sum_{i=1}^{d}\left(\mathbf{\bar{\theta}}_{(i)}^{(l,t)}-\mathbf{\bar{\theta}}_{(i)}^{(l,0)}\right)^{2}\leq\sum_{i=1}^{d}\left(\mathbf{\bar{\theta}}_{i}^{(l,t)}-\mathbf{\bar{\theta}}_{i}^{(l,0)}\right)^{2}$. The Wasserstein regularizer, like variants of L2, regularize the individual parameters and may not preserve the properties of the entire parameter matrix at each layer.
A.7 Details Regarding Spectral Regularization
@@
L2 regularization is not as effective as spectral regularization because it constrains the magnitude of the parameters, and can even cause rank collapse [17]. Spectral regularization constrains only the spectral norm of the weight matrix close to 1, $\sigma^{(1)}(\theta_{l})\approx 1$. This has the effect of controlling the magnitude of the maximal column of the weight matrix. That is, if $\theta\in\mathbb{R}^{d_{l}\times d_{l-1}}$, then $\nicefrac{{1}}{{\sqrt{d_{l}}}}\|\theta_{l}\|_{1}<\|\theta_{l}\|_{2}$ where $\|\theta_{l}\|_{1}=\max_{1\leq j\leq d_{l}}\sum_{i=1}^{d_{l}}|[\theta_{l}]_{ij}|$. This implies that $\sum_{i=1}^{d_{l}}|[\theta_{l}]_{ij}|<\sqrt{d_{l}}$ for every column $j$. This differs in two ways from L2 regularization: (i) the regularizer does not regularize all the columns of the weight matrix, only the parameters in the column with a maximal sum, and (ii) the regularizer does not regularize the individual weights of the column, but their sum. This means that the parameters can potentially be larger, and move further from the initialization, potentially enabling more effective use of the neural network’s capacity. Lastly, we note that L2 regularization can also be viewed in terms of a matrix norm, specifically the Frobenius norm.
@@
The spectral norm can also be controlled via spectral normalization [43]. For continual learning, spectral regularization is preferable over spectral normalization for two reasons: (i) spectral normalization is data-dependent (see Section 2.1 and Equation 12 in [43], and Appendix A.8 for more details), which can be problematic in continual learning due to the changing data distribution, and, (ii) other forms of normalization are often already used to train deep neural networks, such as LayerNorm [44]. LayerNorm, in particular, regularizes the magnitude of the layerwise map by dividing by the norm of the features in the layer. However, because the spectral norm of the underlying parameters still grows at a rate of $\sqrt{t}$, the normalization layer itself will suffer from loss of gradient diversity. For the purposes of ensuring continual trainability, regularization explicitly controls the spectral norm of all parameters and is thus preferable over normalization in isolation. That being said, normalization can still improve performance and optimal performance may be achieved by a combination of regularization and normalization, which we show in Section 5.2.
A.8 Categorizing Regularizers for Continual Learning
We can categorize regularizers broadly into data-dependent and data-indepedent regularizers. Data-dependent regularizers use the data in some way, which can be probelmatic for continual learning when the data distribution changes. Regularizing on one distribution of data does not necessarily maintain useful properties for learning on the next distribution of data. Examples of data-dependent regularizers include feature rank regularization [62], auxiliary tasks and other feature-space regularizers [14]. Other non-standard approaches that have a data-dependent regularization effect include gradient regularizers such as gradient-clipping and weight reinitialization based on dormant neurons [23]. These data-dependent regularizers regularizers control some property on data from the current task but may not control the property on data from a new task, which is needed to maintain plasticity.
Data-Independent Regularization, on the other hand, does not depend on any data. This category regularizes the parameters directly, which is particularly useful in continual learning when the data distribution is changing. Examples of data-independent regularization include L2 regularization towards zero, weight decay (for adaptive gradient methods), and regenerative regularization [17]. Our proposed approach of spectral regularization is data-independent and, moreover, explicitly targets a trainability condition for continual learning. Thus, we expect that it is particularly effective on maintaining trainability in continual learning compared to other data-independent regularizers.
A.9 Spectral Regularization of Other Layers
@@
Other layers using per-unit scaling, such as normalization layers [51, 44], are also spectrally regularized. For example, denote $\gamma\in\mathbb{R}^{d}$ as the trainable scaling parameters for LayerNorm [44], then the element-wise product of the weights can be written as a diagonal weight matrix, $f(z)=\gamma\circ z=\text{Diag}(\gamma)z$. The maximum singular value of a diagonal matrix is the maximum entry, $\sigma_{max}(\text{Diag}(\gamma))=\max_{i}|\gamma_{i}|$. However, optimisation with a maximum is problematic because the max is not differentiable. Furthermore, differentiable surrogates, such as log-sum-exp ($\left(\log\left(\sum_{i}\exp(|\gamma_{i}|)\right)-1\right)^{2}$), only give a loose upper bound on the maximum which is problematic because we want to regularize the maximum towards one. Thus, we regularize each weight towards 1.
@@
Similar to other work, we reshape the convolutional weight tensor (with kernel size ($k\times k$, $d_{in}$ filters and $d_{out}$ filters) into a matrix of size $d_{out}\times(k\cdot k\cdot d_{in})$ [42]. The spectral norm of this reshaped matrix provides an efficient upper bound on the spectral norm of the Toeplitz matrix defining the convolution [63, Corollary 1].
@@ Appendix B Experiment Details
All of our experiments used Adam [64] where the default step size of $0.001$ was selected after an initial sweep over $[0.005,0.001,0.0005]$. For all of our results, we use 10 random seeds and provided a shaded region corresponding to the standard error of the mean. For experiments on SVHN2, CIFAR10 and CIFAR100, we used 5 seeds to sweep over the regularization strengths of $[0.01,0.001,0.0001]$, and found that $0.0001$ worked well on CIFAR10 and CIFAR100 for all regularizers, whereas $0.001$ worked best on SVHN2 for all regularizers.
@@
  • CIFAR10: All of the 50000 images were used for training, 1000 images from the test set were used for validation and the rest were used for testing. We used a batch size of 250 and found this to be effective. For random label non-stationarity, 20 epochs per task and 25 tasks was enough to lose trainability for an unregularized network. For pixel permutation non-stationarity, 10 epochs and 100 tasks was enough to lose trainability for an unregularized network.
  • CIFAR100: All of the 50000 images were used for training, 1000 images from the test set were used for validation and the rest were used for testing. We used a batch size of 250 and found this to be effective. For random label non-stationarity, 20 epochs per task and 25 tasks was enough to lose trainability for an unregularized network. For pixel permutation non-stationarity, 10 epochs and 100 tasks was enough to lose trainability for an unregularized network.
@@
  • For MNIST, EMNIST, and Fashion MNIST: 4-layer MLP with 256 neurons per layer and relu activations. For experiments that use it, LayerNorm was applied after the linear weight matrix and before the non-linearity.
  • For SVHN2, CIFAR10, and CIFAR100: An off-the-shelf ResNet-18 with batch norm, as well as an off-the-shelf Vision Transformer (tiny).

### v2

@@ 6 Conclusion
In this paper, we investigated the connection between initialization, trainability, and regularization. We identified that deviations of the maximum singular value of each layer can lead to low gradient diversity, preventing neural networks from training on new tasks. To directly preserve trainability properties present at initialization, we proposed spectral regularization as a way of controlling the maximum singular value of each layer so that it does not deviate significantly from $1$. Our experiments show that spectral regularization is more performant and less sensitive to hyperparameters than other methods across datasets, nonstationarities, and architectures. While several methods mitigate loss of trainability to a varying degree, learning continually with spectral regularization is robust, achieving high initial and sustained performance. We also showed that spectral regularization is capable of improving generalization with both Vision Transformer and ResNet-18 on continual versions of tiny-ImageNet, CIFAR10, CIFAR100, amongst others. Lastly, we showed that spectral regularization learned a more effective policy in reinforcement learning by avoiding early loss of plasticity due to primacy bias.
@@ Appendix A Additional Details
A.1 Illustrative Example: Large Spectral Norm Can Impede Trainability
We show that, given a family of solutions on the first task, preference should be given to solutions that have a spectral norm that is closer to one. We consider 2 dimensional binary and orthogonal inputs, $x_{1}=[0,1]^{\top}$ and $x_{2}=[1,0]^{\top}$, with corresponding binary targets, $y_{1}=y_{2}=1$. We consider the squared loss, and use a mutli-layer perceptron with one hidden layer and two units, $f_{w_{2},w_{1}}(x)=w_{2}\texttt{ReLU}(w_{1}x)$. A common observation regarding solutions found for deep neural networks is that they are low rank (Arora et al., 2019, Jacot, 2023, Timor et al., 2023, Galanti et al., 2022). In addition, it has been observed that large outliers in the spectrum of a deep network emerge during training, meaning that the largest singular values of the weight matrices tend to be much larger than their smallest singular values (Pennington et al., 2018, Martin and Mahoney, 2021). Consider one such solution where $w_{1}=\begin{bmatrix}-1&c\\ \nicefrac{{1}}{{c}}&-1\end{bmatrix}$, and $w_{2}=[\nicefrac{{1}}{{c}},c]\in\mathbb{R}^{1\times 2}$. The input-output function is invariant under different values of $c$, achieving zero error for any particular choice. However, the parameter matrix for the hidden activations, $w_{1}$, is low rank because it only has a single non-zero singular value that is not invariant. That is, the largest singular value of $w_{1}$ depends on the choice of $c$, $\|w_{1}\|_{2}:=\sigma_{max}(w_{1})=c+\nicefrac{{1}}{{c}}$. In addition, the gradients and parameter updates depend on $c$. Thus, if the targets were to change, $y_{1}^{\prime}=0$, the imbalance of singular values due to very large or small values of $c$ can impede trainability:
@@ A.1 Illustrative Example: Large Spectral Norm Can Impede Trainability
where $\ell_{1}=\ell(f_{w_{2},w_{1}}(x_{1}),y^{\prime}_{1})$, and $\alpha$ is a step-size. One of the consequences of the imbalanced singular values is that the per-example gradients are poorly conditioned (Wu et al., 2021). That is, when we sample $x_{1}$, there will be a large gradient for $w_{2}$, $\|\nabla_{w_{2}}\ell(f_{w_{2},w_{1}}(x_{1}),y_{1}^{\prime})\|^{2}=c$ but a small gradient for $w_{1}$, $\|\nabla_{w_{1}}\ell(f_{w_{2},w_{1}}(x_{1}),y_{1}^{\prime})\|^{2}=\nicefrac{{1}}{{c}}$. Thus, a sufficiently small step-size that stabilizes the update to $w_{2}$ leads to slow learning on $w_{1}$. A similar result holds if the target $y_{2}$ were to change In Figure 6, we demonstrate that learning is more stable with a smaller spectral norm in this example. For higher dimensional weight matrices, we may expect that the imbalance of the singular values leads to low gradient diversity where particular per-example gradients give rise to outliers with a large gradient magnitudes. Thus, all else being equal, a solution with a smaller spectral norm ($c\approx 1$) is preferable for the purposes of continual learning.
@@
A.2 Differences Between Continual and Non-continual Learning
@@
A.3 Spectral Norm and Lipschitz Constant
One other potential reason for why a growing spectral norm is problematic is from the theory convex optimisation. For the ReLU activation function, the Lipschitz constant of the layer map, $h_{l+1}=\phi(\theta_{l}h_{l})$, is equal to the spectral norm of the weight matrix. We can thus bound the Lipschitz constant of the entire neural network above by the product of the layer-wise Lipschitz constants (Szegedy et al., 2013). Putting these results together, we conclude that during continual learning, the Lipschitz constant of each layer is increasing and thus the Lipschitz constant of the entire network is also increasing. This is problematic from the standpoint of optimisation because gradient descent only converges on convex optimisation problems when the step-size is smaller than $\frac{2}{L}$, where $L$ is the Lipschitz constant. Thus, if the Lipschitz constant grows then, eventually, the chosen step-size will be too large and gradient descent will not converge locally on the task.
A.4 More Details on Regularization and Trainability
One other important advantage of regularization is that it has well understood effects for ensuring trainability in approaches outside of deep learning, such as degenerate linear regression and inverse problems (Benning and Burger, 2018).
@@ A.4 More Details on Regularization and Trainability
A.5 Details Regarding Spectral Regularization
@@
L2 regularization is not as effective as spectral regularization because it constrains the magnitude of the parameters, and can even cause rank collapse (Kumar et al., 2023). Spectral regularization constrains only the spectral norm of the weight matrix close to 1, $\sigma^{(1)}(\theta_{l})\approx 1$. This has the effect of controlling the magnitude of the maximal column of the weight matrix. That is, if $\theta\in\mathbb{R}^{d_{l}\times d_{l-1}}$, then $\nicefrac{{1}}{{\sqrt{d_{l}}}}\|\theta_{l}\|_{1}<\|\theta_{l}\|_{2}$ where $\|\theta_{l}\|_{1}=\max_{1\leq j\leq d_{l}}\sum_{i=1}^{d_{l}}|[\theta_{l}]_{ij}|$. This implies that $\sum_{i=1}^{d_{l}}|[\theta_{l}]_{ij}|<\sqrt{d_{l}}$ for every column $j$. This differs in two ways from L2 regularization: (i) the regularizer does not regularize all the columns of the weight matrix, only the parameters in the column with a maximal sum, and (ii) the regularizer does not regularize the individual weights of the column, but their sum. This means that the parameters can potentially be larger, and move further from the initialization, potentially enabling more effective use of the neural network’s capacity. Lastly, we note that L2 regularization can also be viewed in terms of a matrix norm, specifically the Frobenius norm.
@@
The spectral norm can also be controlled via spectral normalization (Miyato et al., 2018). For continual learning, spectral regularization is preferable over spectral normalization for two reasons: (i) spectral normalization is data-dependent (see Section 2.1 and Equation 12 in Miyato et al. (2018), and Appendix A.6 for more details), which can be problematic in continual learning due to the changing data distribution, and, (ii) other forms of normalization are often already used to train deep neural networks, such as LayerNorm (Ba et al., 2016). LayerNorm, in particular, regularizes the magnitude of the layerwise map by dividing by the norm of the features in the layer. However, because the spectral norm of the underlying parameters still grows at a rate of $\sqrt{t}$, the normalization layer itself will suffer from loss of gradient diversity. For the purposes of ensuring continual trainability, regularization explicitly controls the spectral norm of all parameters and is thus preferable over normalization in isolation. That being said, normalization can still improve performance and optimal performance may be achieved by a combination of regularization and normalization, which we show in Section 5.3.
A.6 Categorizing Regularizers for Continual Learning
We can categorize regularizers broadly into data-dependent and data-indepedent regularizers. Data-dependent regularizers use the data in some way, which can be probelmatic for continual learning when the data distribution changes. Regularizing on one distribution of data does not necessarily maintain useful properties for learning on the next distribution of data. Examples of data-dependent regularizers include feature rank regularization (Kumar et al., 2021), auxiliary tasks and other feature-space regularizers (Lyle et al., 2022). Other non-standard approaches that have a data-dependent regularization effect include gradient regularizers such as gradient-clipping and weight reinitialization based on dormant neurons (Sokar et al., 2023). These data-dependent regularizers regularizers control some property on data from the current task but may not control the property on data from a new task, which is needed to maintain plasticity.
Data-Independent Regularization, on the other hand, does not depend on any data. This category regularizes the parameters directly, which is particularly useful in continual learning when the data distribution is changing. Examples of data-independent regularization include L2 regularization towards zero, weight decay (for adaptive gradient methods), and regenerative regularization (Kumar et al., 2023). Our proposed approach of spectral regularization is data-independent and, moreover, explicitly targets a trainability condition for continual learning. Thus, we expect that it is particularly effective on maintaining trainability in continual learning compared to other data-independent regularizers.
A.7 Spectral Regularization of Other Layers
@@
Other layers using per-unit scaling, such as normalization layers (Ioffe and Szegedy, 2015, Ba et al., 2016), are also spectrally regularized. For example, denote $\gamma\in\mathbb{R}^{d}$ as the trainable scaling parameters for LayerNorm (Ba et al., 2016), then the element-wise product of the weights can be written as a diagonal weight matrix, $f(z)=\gamma\circ z=\text{Diag}(\gamma)z$. The maximum singular value of a diagonal matrix is the maximum entry, $\sigma_{max}(\text{Diag}(\gamma))=\max_{i}|\gamma_{i}|$. However, optimisation with a maximum is problematic because the max is not differentiable. Furthermore, differentiable surrogates, such as log-sum-exp ($\left(\log\left(\sum_{i}\exp(|\gamma_{i}|)\right)-1\right)^{2}$), only give a loose upper bound on the maximum which is problematic because we want to regularize the maximum towards one. Thus, we regularize each weight towards 1.
@@
Similar to other work, we reshape the convolutional weight tensor (with kernel size ($k\times k$, $d_{in}$ filters and $d_{out}$ filters) into a matrix of size $d_{out}\times(k\cdot k\cdot d_{in})$ (Yoshida and Miyato, 2017). The spectral norm of this reshaped matrix provides an efficient upper bound on the spectral norm of the Toeplitz matrix defining the convolution (Tsuzuku et al., 2018, Corollary 1).
@@ Appendix B Experiment Details
All of our experiments used Adam (Kingma and Ba, 2015) where the default step size of $0.001$ was selected after an initial sweep over $[0.005,0.001,0.0005]$. For all of our results, we use 10 random seeds and provided a shaded region corresponding to the standard error of the mean. For experiments on tiny-ImageNet SVHN2, CIFAR10 and CIFAR100, we used 4 seeds to sweep over the regularization strengths of $[0.01,0.001,0.0001]$, and found that $0.0001$ worked well on tiny-ImageNet, CIFAR10 and CIFAR100 for all regularizers, whereas $0.001$ worked best on SVHN2 for all regularizers.
@@
  • CIFAR10: All of the 50000 images were used for training, 1000 images from the test set were used for validation and the rest were used for testing. We used a batch size of 250 and found this to be effective. For random label non-stationarity, 20 epochs per task and 30 tasks was enough to lose trainability for an unregularized network. For pixel permutation non-stationarity, 10 epochs and 100 tasks was enough to lose trainability for an unregularized network.
  • CIFAR100: All of the 50000 images were used for training, 1000 images from the test set were used for validation and the rest were used for testing. We used a batch size of 250 and found this to be effective. For random label non-stationarity, 20 epochs per task and 30 tasks was enough to lose trainability for an unregularized network. For pixel permutation non-stationarity, 10 epochs and 100 tasks was enough to lose trainability for an unregularized network.
  • tiny-ImageNet: All of the 100000 images were used for training, 10000 images were used for validation, and 10000 images were used for testing according to the predetermined split. We used a batch size of 250 and found this to be effective. For random label non-stationarity, 20 epochs per task and 30 tasks was enough to lose trainability for an unregularized network. For pixel permutation non-stationarity, 20 epochs and 100 tasks was enough to lose trainability for an unregularized network.
  • Reinforcement Learning: We evaluate spectral regularization in RL control tasks from DMC benchmark (Tassa et al., 2020), with SAC method (Haarnoja et al., 2018) in demanding replay ratio (RR) regime (D’Oro et al., 2022) with 16 gradient updates per every new environments step. We choose this setup because a high RR regime leads to significant primacy bias (Nikishin et al., 2022) defined as a tendency to overfit initial experiences that damages the rest of the learning process.
@@
  • MNIST, EMNIST, and Fashion MNIST: 4-layer MLP with 256 neurons per layer and relu activations. For experiments that use it, LayerNorm was applied after the linear weight matrix and before the non-linearity.
  • tiny-ImageNet CIFAR10, CIFAR100, and SVHN2: An off-the-shelf ResNet-18 with batch norm, as well as an off-the-shelf Vision Transformer (tiny).
  • Reinforcement Learning: Recent studies (Nauman et al., 2024, Ball et al., 2023) have demonstrated that, in this setup, only resets with layer normalization (Ba et al., 2016) serve as an effective mitigation strategy. We compare the performance of the SAC agent with both spectral regularization and layer normalization to two baseline agents: SAC with only layer normalization, and SAC with layer normalization plus resets. We use spectral regularization with a coefficient $41e-$ 4 for both the actor and critic. For every method, we use a single critic and architecture size of $2$ layers and $256$ neurons per layer for both actor and critic. Using a random policy, we prefill a replay buffer with $10,000$ transitions before starting the training. Replay buffer maximum size is $1$ million transitions.